# A Survey of Deep Learning for Complex Speech Spectrograms

### v1

@@ 6.3 Speech Separation
Since the introduction of deep cluster (Hershey et al. (2016)) and permutation invariant training (PIT) (Yu et al. (2017); Kolbæk et al. (2017)), the performance of DNN-based speech separation methods has seen substantial improvements. Early approaches incorporated phase information using the phase sensitive mask (PSM) (Erdogan et al. (2015)), which yielded promising results in speech separation (Kolbæk et al. (2017)). However, PSM is applied only to the magnitude spectrogram, while the mixture phase is still used for inverse STFT, limiting its effectiveness. A more suitable approach involves estimating the real and imaginary spectrograms of separated sources. In this direction, Williamson et al. (2016) proposed using the complex-valued ideal ratio mask (cIRM) for speech separation. Both objective and subjective evaluations demonstrated the effectiveness of estimating phase spectrogram in improving separation quality. Additionally, Wang et al. (2021a) introduced a direct complex spectrogram mapping approach, training the model with a novel multi-domain loss. Compared to conventional objective functions, this novel loss function led to significant performance improvements in speech separation (Wang et al. (2021b)). Despite these advancements, most speech separation research has focused on time-domain approaches, which require smaller window and hop sizes and have demonstrated superior performance. Most recently, Wang et al. (2023) proposed the TF-GridNet framework, which leverages both global and local information through a multi-path model. This model integrates intra-frame, sub-band and full-band modules to enhance separation performance. Experimental results indicate that TF-GridNet outperforms other state-of-the-art time-domain speech separation baselines on the SI-SDR metric.
@@
It is worth noting that this subsection offers only a concise overview of monaural speech separation, with an emphasis on complex spectrogram processing. For a more comprehensive review of speech separation techniques, readers are referred to the recent work by Araki et al. (2025).
@@ 7 Generative Models for Complex Spectrograms
GANs (Goodfellow et al. (2014)) represent a class of powerful generative models, composed of two neural networks — a generator and a discriminator — trained in a minimax game. While classical GANs learn to generate realistic data from random noise in an unsupervised manner (Goodfellow et al. (2014)), recent applications in complex spectrogram processing typically adopt the framework of conditional GANs, where generation is guided by an additional input condition. Applications of GANs on complex spectrograms processing are mostly focusing on speech enhancement, and also phase retrieval and audio synthesis.
In speech enhancement, the condition is the noisy speech signal itself, which implicitly contains both noise and the underlying clean speech. From a generative modeling perspective, the objective is to learn to conditionally generate clean complex spectrograms from noisy inputs. Specifically, the generator is trained to produce enhanced spectrograms using either real-valued or complex-valued neural networks (Kothapally and Hansen (2022); Lu et al. (2023); Kühne et al. (2025); Cao et al. (2022)). The discriminator, meanwhile, plays a critical role in shaping the learning dynamics by evaluating how realistic the generated spectrograms are. Beyond basic real-vs-fake classification (Kothapally and Hansen (2022)), many studies have enhanced the discriminator with perceptual constraints by introducing feature-level distance metrics or differentiable perceptual evaluation scores, such as PESQ (Cao et al. (2022); Zhang et al. (2024); Lu et al. (2023)). These additions enable the model to better align with human auditory perception, reinforcing the idea of GANs as perceptually-aware generative models.
@@
Diffusion models (Song et al. (2021)) are generative models that learn to gradually denoise data by reversing a fixed forward noising process, typically formulated as a Markov chain or a continuous stochastic differential equation (SDE). The works of applying diffusion models to complex spectrograms are predominantly in speech enhancement. Similar with the cases in GAN, diffusion models can be regarded as conditional generative models in such scenarios. Building upon the foundation established by Song et al. (2021), Welker et al. (2022) introduced a continuous SDE framework for speech enhancement, named SGMSE. This generalized continuous formulation enables flexible integration of arbitrary numerical solvers during the reverse diffusion process. As the model is proposed for complex spectrograms, the perturbation kernel used in SGMSE is a multivariant circularly-symmetric complex Gaussian distribution $\mathcal{N}_{C}(\mu,\sigma)$. Moreover, advancements of diffusion models have been further used to handle complex spectrogram-based enhancement, including cold diffusion (Yen et al. (2023)), unconditional diffusion model (Lemercier et al. (2023b)). Recently, Gonzalez et al. (2024) extends a diffusion model originally proposed for image generation to incorporate a speech enhancement system, where the diffusion process has a non-zero long-term mean equal to the conditioning signal. This approach directly models the clean speech signal and offers a more general formulation. Complex-valued backbones like DCUNET (Choi et al. (2019)) and DCCRN (Hu et al. (2020)), real-valued backbones include Noise Conditional Score Network (NCSN++) (Song et al. (2021)), CRN (Tan and Wang (2019b)) have been explored for diffusion model-based complex spectrogram processing across works (Welker et al. (2022); Yen et al. (2023); Richter et al. (2023); Wang and Wang (2023)). For a more comprehensive exploration of this topic, Lemercier et al. (2024) offered a detailed analysis in their work.
@@ 8 Conclusion
This survey has comprehensively explored the advancements and methodologies in processing complex spectrograms using DNNs. We delved into the fundamental aspects of complex spectrograms and the key components of complex-valued neural networks. Various training strategies and loss functions were discussed for learning complex spectrograms. Additionally, we examined the DNN-based applications in phase retrieval, speech enhancement, and speech separation, showcasing their effectiveness in improving audio signal quality. Finally, the intersection of complex spectrograms with generative models was explored.

### v2

@@ 7.3 Speaker Separation
Since the introduction of deep cluster (Hershey et al. (2016)) and permutation invariant training (PIT) (Yu et al. (2017); Kolbæk et al. (2017)), the performance of DNN-based speaker separation methods has seen substantial improvements. Early approaches incorporated phase information using the phase sensitive mask (PSM) (Erdogan et al. (2015)), which yielded promising results in speaker separation (Kolbæk et al. (2017)). However, PSM is applied only to the magnitude spectrogram, while the mixture phase is still used for inverse STFT, limiting its effectiveness. A more suitable approach involves estimating the real and imaginary spectrograms of separated sources. In this direction, Williamson et al. (2016) proposed using the complex-valued ideal ratio mask (cIRM) for speaker separation. Both objective and subjective evaluations demonstrated the effectiveness of estimating phase spectrogram in improving separation quality. Additionally, Wang et al. (2021b) introduced a direct complex spectrogram mapping approach, training the model with a novel multi-domain loss. Compared to conventional objective functions, this novel loss function led to significant performance improvements in speaker separation (Wang et al. (2021c)). Despite these advancements, most speaker separation research has focused on time-domain approaches, which require smaller window and hop sizes and have demonstrated superior performance. Recently, Wang et al. (2023a) proposed the TF-GridNet framework, which leverages both global and local information through a multi-path architecture. This model integrates intra-frame, sub-band, and full-band modules to enhance separation performance. Experimental results demonstrate that TF-GridNet outperforms other state-of-the-art time-domain speaker separation baselines in terms of the SI-SDR metric. Subsequently, SpatialNet (Quan and Li (2024)) and TF-CrossNet (Kalkhorani and Wang (2024)) have been introduced, achieving superior results with reduced computational cost in speaker separation.
@@
It is worth noting that this subsection offers only a concise overview of monaural speaker separation, with an emphasis on complex spectrogram processing. For a more comprehensive review of speaker separation techniques, readers are referred to the recent work by Araki et al. (2025) and Li et al. (2025).
@@ 8 Generative Models for Complex Spectrograms
GANs (Goodfellow et al. (2014)) represent a class of powerful generative models, composed of two neural networks — a generator and a discriminator — trained in a minimax game. While classical GANs learn to generate realistic data from random noise in an unsupervised manner (Goodfellow et al. (2014)), recent applications in complex spectrogram processing typically adopt the framework of conditional GANs, where generation is guided by an additional input condition. Applications of GANs on complex spectrograms processing are mostly focusing on speech enhancement, and also phase retrieval and audio synthesis. In speech enhancement, the condition is the noisy speech signal itself, which implicitly contains both noise and the underlying clean speech. From a generative modeling perspective, the objective is to learn to conditionally generate clean complex spectrograms from noisy inputs. Specifically, the generator is trained to produce enhanced spectrograms using either real-valued or complex-valued neural networks (Kothapally and Hansen (2022); Lu et al. (2023); Kühne et al. (2025b); Cao et al. (2022)). The discriminator, meanwhile, plays a critical role in shaping the learning dynamics by evaluating how realistic the generated spectrograms are. Beyond basic real-vs-fake classification (Kothapally and Hansen (2022)), many studies have enhanced the discriminator with perceptual constraints by introducing feature-level distance metrics or differentiable perceptual evaluation scores, such as PESQ (Cao et al. (2022); Zhang et al. (2024a); Lu et al. (2023)). These additions enable the model to better align with human auditory perception, reinforcing the idea of GANs as perceptually-aware generative models.
@@
Diffusion models (Song et al. (2021)) are generative models that learn to gradually denoise data by reversing a fixed forward noising process, typically formulated as a Markov chain or a continuous stochastic differential equation (SDE).
The works of applying diffusion models to complex spectrograms are predominantly in speech enhancement. Similar with the cases in GAN, diffusion models can be regarded as conditional generative models in such scenarios. Building upon the foundation established by Song et al. (2021), Welker et al. (2022) introduced a continuous SDE framework for speech enhancement, named SGMSE. This generalized continuous formulation enables flexible integration of arbitrary numerical solvers during the reverse diffusion process. As the model is proposed for complex spectrograms, the perturbation kernel used in SGMSE is a multivariant circularly-symmetric complex Gaussian distribution $\mathcal{N}_{C}(\mu,\sigma)$. Moreover, advancements of diffusion models have been further used to handle complex spectrogram-based enhancement, including cold diffusion (Yen et al. (2023)), unconditional diffusion model (Lemercier et al. (2023b)). Recently, Gonzalez et al. (2024) extends a diffusion model originally proposed for image generation to incorporate a speech enhancement system, where the diffusion process has a non-zero long-term mean equal to the conditioning signal. This approach directly models the clean speech signal and offers a more general formulation. Complex-valued backbones like DCUNET (Choi et al. (2019)) and DCCRN (Hu et al. (2020)), real-valued backbones include Noise Conditional Score Network (NCSN++) (Song et al. (2021)), CRN (Tan and Wang (2019b)) have been explored for diffusion model-based complex spectrogram processing across works (Welker et al. (2022); Yen et al. (2023); Richter et al. (2023); Wang and Wang (2023)).
Additionally, diffusion models are often combined with discriminative models in a hybrid framework. The discriminative model first provides an initial estimate of clean speech, which is then refined by the diffusion model to further reduce distortions using techniques such as denoising diffusion restoration and stochastic regeneration (Wang and Wang (2023); Sawata et al. (2023); Shi et al. (2024); Lemercier et al. (2023a)). This hybrid method avoids generating artifacts such as vocalizing and breathing effects while significantly reducing computational costs (Lemercier et al. (2023a)). More recently, low-latency diffusion model-based speech enhancement has been explored (Lay et al. (2025)). For a more comprehensive exploration of this topic, Lemercier et al. (2024) offered a detailed analysis in their work.
@@ 9 Conclusion
This survey has comprehensively explored the advancements and methodologies in processing complex spectrograms using DNNs. We delved into the fundamental aspects of complex spectrograms and the key components of CVNNs. The approaches of using RVNNs for handling complex spectrograms have also been summarized. Various training strategies and loss functions were discussed for learning complex spectrograms. Additionally, we examined the DNN-based applications in phase retrieval, speech enhancement, and speaker separation, showcasing their effectiveness in improving audio signal quality. Finally, the intersection of complex spectrograms with generative models was explored.